OpenStudy (anonymous):

evaluate the integral and interpret it as the area of a region. sketch the region.

OpenStudy (anonymous):

\[\int\limits_{0}^{\Pi/2}\] absolute value (sinx- cosx) dx

OpenStudy (anonymous):

 it is twice the integral from zero to pi/2 of (cosx-sinx)

OpenStudy (anonymous):

? i still dont understand. i evaluated it and got 1 as an answer but its not the right answer

OpenStudy (anonymous):

sorry i meant twice the intergral from zero to pi/4. what is the answer?

OpenStudy (anonymous):

\[(3/2)\sqrt{3}-1\] thats the answer

OpenStudy (anonymous):

well i don't know if that can be the correct answer. you have to consider the graph in two parts. from zero to pi/4 the function is negative since sin(x) is less than cos(x) so you have to intergrate the function as cos(x)-sin(x). from pi/4 to pi/2 the function is positive so you have to intergrate it as sin(x)-cos(x). however the two areas are the same so you could just multiply the area of the intergral from 0 to pi/4 of (cos(x)-sin(x))

OpenStudy (amistre64):

\[\int |x|`dx=\frac12x|x|+C\]

OpenStudy (amistre64):

say: u = sinx -cosx du = cosx - sinx dx\[-\int\frac{|u|}{u}du\]

OpenStudy (amistre64):

which works out to be:\[-\frac{u^2}{|u|}\]now when x=0; u=sin(0)-cos(0)=-1 when x=pi/2; u=sin(pi/2)-cos(pi/2)=1 \[-\left.\frac{u^2}{|u|}\right|_{-1}^{1}=-1+1=0\] if i did it right

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=integrate+%7Csinx-cosx%7C+from+0+to+pi%2F2 the wolf says i did it wrong :)